OpenStudy (anonymous):

Find the dimensions of a triangle if the base is (2/3) the measure of the height and the area is 12 square centimeters.

OpenStudy (anonymous):

ohhh I like your avatar. Nyan cat! Back on topic.... Area of triangle is A=(1/2b) *h

OpenStudy (anonymous):

my bad. A= 1/2*b*h

OpenStudy (anonymous):

so plug in 12 for A.

OpenStudy (anonymous):

base = 2/3 * height Area = 1/2 * base * height Area = 1/2 * 2/3 * height * height = 12 height^2 = 36 height = 6 base = 4

OpenStudy (cathyangs):

ok. The formula for area of a triangle is (1/2)*base*height. if b [base] is 2/3 of h [height], then we can write b=(2/3)h so then the area, 12, is equal to (1/2)(2/3)h*h which is (2/6)h^2, or (1/3)h^2 and, doing some algebra, the height is 6, so the base is 4.

OpenStudy (anonymous):

Step 1. A=1/2*b*h Step 2. 12=1/2*b*h Step 3. 24=b*h (muliply by 2 on both sides) Step 4. You know that base is 2/3 of height so you can express as: b = 2/3* h. Plug this in. Step 5. 24=2/3*h*h Step 6. 36= h ^2 Step 7. h = 6

OpenStudy (anonymous):

in step 6 I just mulitplied both sides by 3/2

OpenStudy (anonymous):

No medals. No thanks.

OpenStudy (anonymous):

i found the area and the base but it says to find the dimensions so am i done?

OpenStudy (anonymous):

and height

OpenStudy (cathyangs):

yes! :D

OpenStudy (anonymous):

dimensions means the base, the height, and area is already given :]. but you also need the 3rd side of the triangle.

OpenStudy (anonymous):

Dimensions in this case means both base and height. 3 people showed you how to get both.

OpenStudy (cathyangs):

In this case, I think the dimensions are base and height.

OpenStudy (anonymous):

just use 3,4,5. The special right triangle...I think you can use it in this case...not sure. Been a while since I had math

OpenStudy (anonymous):

ok ty everyone TACTICAL MEDAL INBOUND

OpenStudy (anonymous):

or you can use pythag's theorem. a^2=b^2+c^2

OpenStudy (anonymous):

oh shoot don't listen to that last comment it is a^2+b^2=c^2

OpenStudy (anonymous):

silly me

OpenStudy (anonymous):

Please don't forget the 3rd side of the triangle! A triangle has 3 sides!!! base and height won't be sufficient dimensions